### v1

# Characterization of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements

Abstract
Future gravitational wave detectors are going to cool down their test masses to cryogenic temperatures to reduce thermal noise. Crystalline materials are considered most promising for making test masses and their coatings due to their excellent thermal and optical properties at low temperatures. However, birefringence due to local impurities and inhomogeneities in the crystal can degrade the performance of the detector. Birefringence measurement, or birefringence mapping over a two-dimensional area has become important. This paper describes a method of fast birefringence measurement for a large sample by simply combining a series of transmission wavefront error measurements using linearly polarized light with Fizeau interferometers. With this method, the birefringence inhomogeneity of two KAGRA’s input test masses with diameter of 22 cm is fully reconstructed. The birefringence information is then used to calculate the transverse beam shape of the light fields in orthogonal polarization directions when passing through the substrate. The calculated beam shapes show a good match with in-situ measurements using KAGRA interferometer. This technique is crucial for birefringence characterization of test masses in future detectors where even larger size are used.
I Introduction
Thermal noise is one of the main noise sources for future gravitational wave detectors [1, 2]. To reduce thermal noise, future detectors like LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures. Cosmic Explorer [6] also considers cryogenic operation for an upgrade. KAGRA [7, 8, 9, 10] is the only one operating at cryogenic temperatures and pioneering this technique among current large scale ground-based detectors. All these detectors apply large size high-quality mirrors as their test masses forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves. Fused silica is commonly used to make test masses at room temperature, but it is not compatible with low temperatures. Crystalline materials are considered most promising to make cryogenic test masses. Sapphire was selected as the material for KAGRA’s test masses [11] due to its excellent thermal and optical properties at low temperature [12, 13]. However, recent studies reveal that birefringence of crystalline substrates and coatings can degrade the performance of the detector [14, 15, 16, 17, 18, 19]. In this case, the birefringence fluctuates in magnitude and direction throughout the substrate of the mirror. Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on the optical performance.
There are many studies on two-dimensional birefringence measurements for transparent materials [14, 15, 20, 21, 22, 23, 24]. These measurements are either using a single-beam scanning technique, or expanding the beam to a large aperture and imaging with a sensor array. The scanning method is rather time-consuming for large size samples, e.g., mirrors with diameters usually more than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [25, 26]. The phase-shifting method in digital interferometry has been previously proposed for fast 2D birefringence measurements with a very high accuracy [27, 28, 29]. Fizeau interferometers are one of the most accurate methods to extract the phase information of an optical component (reflection phase or transmission phase) over a large area. Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test. In the presence of birefringence, however, a simple TWE measurement does not reveal the entire information about the substrate.
In this paper we describe a method of combining a series of TWE map measurements using linearly polarized light in an appropriate way to fully reconstruct the refractive index information including birefringence. With this method, the birefringence inhomogeneity of two KAGRA’s input test masses (ITM) is extracted. This birefringence information is then used to calculate the transverse beam shape of the light fields in orthogonal polarization directions when passing through the substrate. The calculated beam shapes show a good match with in-situ beam shape measurements. The method of extracting birefringence information from TWE measurements with Fizeau interferometers is introduced in Section II. Birefringence characterization of KAGRA’s current ITMs is described in Section III. Our conclusions are summarized in Section IV.
II Generating a birefringence map with a Fizeau interferometer
Figure 1 shows a schematic of a Fizeau interferometer. We analyze the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD. By comparing the results with and without the sample inserted into the cavity formed by the reference planes 1 and 2, we can obtain the information about the optical thickness of the sample. Since the optical thickness is the physical thickness times the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflecting coating) and S2 (surface with highly-reflective coating) precisely and subtracting out the variation of physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
Figure 1: Schematic of TWE measurement with a Fizeau interferometer. S1 is the surface with anti-reflecting coating and S2 is the surface with highly-reflective coating. The interference between light fields $E_{\mathrm{1}}$ and $E_{\mathrm{2}}$ reflected by reference planes 1 and 2 respectively gives the TWE information of the sample.
In order to extract the birefringence information, measurements have to be performed using a linearly polarized light. To understand the behavior of such measurements, let us first consider what happens to a linearly polarized light going through a birefringent substrate.
II.1 Transmission of a linearly polarized light through a birefringent medium
Let us concentrate on an optical axis going through a single column of the substrate as shown in Figure 2. We define our preferred polarization axes called s and p in the vertical and horizontal directions with respect to the laboratory frame. The extra-ordinary and ordinary axes of the crystal in that particular column is assumed to be rotated by an angle $\theta$ with respect to our polarization axes. The input beam to the Fizeau interferometer is purely polarized in s ($E_{\mathrm{s}}$).
Figure 2: Transmission of a linearly polarized optical field through a column of a sample substrate.
We will use the Jones matrix formalism to analyze the conversion of the polarization state [30]. Here, the polarization state of light is represented by a vector,
$\vec{V}=\begin{pmatrix}E_{\mathrm{s}}\\ E_{\mathrm{p}}\end{pmatrix},$ (1)
where $E_{\mathrm{s}}$ and $E_{\mathrm{p}}$ are the complex amplitudes of s- and p-polarized electric fields.
After passing through the substrate twice (a round trip through the substrate), the polarization state is converted to
$\vec{V}^{\prime}=M\cdot\vec{V},\quad\mathrm{with}\quad M=e^{i\alpha_{+}}\begin{pmatrix}e^{i\alpha_{-}}\cos^{2}\theta+e^{-i\alpha_{-}}\sin^{2}\theta&-i\sin\alpha_{-}\sin 2\theta\\ -i\sin\alpha_{-}\sin 2\theta&e^{-i\alpha_{-}}\cos^{2}\theta+e^{i\alpha_{-}}\sin^{2}\theta\end{pmatrix},$ (2)
where $M$ is the round-trip Jones matrix of the substrate. We denote the indices of refraction for the extra-ordinary and ordinary axes as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively. Then we can compute the round trip phase changes of optical fields polarized in those axes:
$\alpha_{\mathrm{e}}\equiv 2\pi\frac{2dn_{\mathrm{e}}}{\lambda},$ (3)
$\alpha_{\mathrm{o}}\equiv 2\pi\frac{2dn_{\mathrm{o}}}{\lambda},$ (4)
where $d$ is the thickness of the substrate, $\lambda$ is the wavelength of the light.
Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes,
$\alpha_{+}\equiv\frac{\alpha_{\mathrm{e}}+\alpha_{\mathrm{o}}}{2},$ (5)
$\alpha_{-}\equiv\frac{\alpha_{\mathrm{e}}-\alpha_{\mathrm{o}}}{2}.$ (6)
In most cases, the common phase is not important and we will ignore $\alpha_{+}$ in the following discussion. Note that the Jones matrix also works for a non-uniform medium (discussed in Appendix A).
For a purely s-polarized input field, $\vec{V}=(1,0)^{\mathrm{T}}$, the field returning from the substrate becomes,
$\vec{V}^{\prime}=\begin{pmatrix}E_{\mathrm{s}}\\ E_{\mathrm{p}}\end{pmatrix}=\begin{pmatrix}e^{i\alpha_{-}}\cos^{2}\theta+e^{-i\alpha_{-}}\sin^{2}\theta\\ -i\sin\alpha_{-}\sin 2\theta\end{pmatrix}.$ (7)
This $\vec{V}^{\prime}$ is the same as the field $E_{2}$ in Figure 1 at the CCD, except for the common phase term, which is ignored.
II.2 Interference with the reference field
The field $\vec{V}^{\prime}$ computed above is superposed with the reference field ($E_{1}$ in Figure 1). The reference field $\vec{V}_{\mathrm{r}}$ is purely polarized in s and has a certain phase offset $\phi$ from $\vec{V}^{\prime}$,
$\vec{V}_{\mathrm{r}}=\begin{pmatrix}E_{\mathrm{r}}\\ 0\end{pmatrix}=\begin{pmatrix}e^{i\phi}\\ 0\end{pmatrix}.$ (8)
The optical power detected by the CCD after the interference of $\vec{V}^{\prime}$ and $\vec{V}_{\mathrm{r}}$ is,
$P=(\vec{V}_{\mathrm{r}}+\vec{V}^{\prime})\cdot(\vec{V}_{\mathrm{r}}+\vec{V}^{\prime})^{*}.$ (9)
By plugging Equations 7 and 8 into Equation 9, we get,
$P$ $=|E_{\mathrm{s}}+E_{\mathrm{r}}|^{2}+|E_{\mathrm{p}}|^{2}$
$=|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}+|E_{\mathrm{r}}|^{2}+E_{\mathrm{s}}E_{\mathrm{r}}^{*}+E_{\mathrm{s}}^{*}E_{\mathrm{r}}.$ (10)
From the conservation of energy, $|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}=1$. From Equation 8, $|E_{\mathrm{r}}|^{2}=1$. We will ignore those constant terms in the following discussion. Then what is interesting to us is,
$\Delta P$ $=E_{\mathrm{s}}E_{\mathrm{r}}^{*}+E_{\mathrm{s}}^{*}E_{\mathrm{r}}$
$=2\cos^{2}\theta\cos(\phi-\alpha_{-})+2\sin^{2}\theta\cos(\phi+\alpha_{-}).$ (11)
If the s-polarization axis is aligned with the extra-ordinary axis ($\theta=0$), $\Delta P=2\cos(\phi-\alpha_{-})$. Usually, a Fizeau interferometer is equipped with a mechanism to change the position of the reference plane 1, for example, by mounting the reference flat on a piezo stage. Therefore, we can scan the relative phase $\phi$ by known amounts and fit the measured variation of $\Delta P$ to know the value of $\alpha_{-}$. In the general case of $\theta\neq 0$, the Fizeau measurement obtains the phase of the field $E_{\mathrm{s}}$, i.e. $\psi=\arg(E_{\mathrm{s}})$.
II.3 Extraction of $\alpha_{-}$ and $\theta$ from several TWE maps
When the orientation of the extra-ordinary axis is not known, which is almost always the case, we need to take several TWE maps with different orientation of the sample or the measurement beam polarization. Then combine them to extract $\theta$ and $\alpha_{-}$.
As explained in the previous section, a TWE measurement by a Fizeau interferometer yields a transmission phase $\psi$ at each point on the sample surface. Let us assume that we rotate the sample by an angle $\eta$, or equivalently the input polarization orientation by $-\eta$. This is equivalent to changing the orientation of the extra-ordinary axis from $\theta$ to $\theta+\eta$.
Let us denote the measured phase with an $\eta$ rotation of the sample as $\psi(\eta)$. From Equation 7,
$\psi(\eta)$ $=\arg\left[E_{\mathrm{s}}(\theta+\eta)\right]$
$=\arg\left[\cos^{2}(\theta+\eta)e^{i\alpha_{-}}+\sin^{2}(\theta+\eta)e^{-i\alpha_{-}}\right]$
$=\arctan\left[\cos(2\theta+2\eta)\tan\alpha_{-}\right].$ (12)
We can see $\psi(\eta)$ as a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape. Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation 12 to obtain the best fit values of $\theta$ and $\alpha_{-}$.
If $\alpha_{-}$ can be assumed to be very small, which is the case for KAGRA mirrors, we can expand Equation 12 with $\alpha_{-}$, leaving only the $\mathcal{O}(\alpha_{-})$ terms, as
$\psi(\eta)$ $\sim\cos(2\theta+2\eta)\alpha_{-}$
$=\left(\cos(2\theta)\cos(2\eta)-\sin(2\theta)\sin(2\eta)\right)\alpha_{-}.$ (13)
Assuming that we have four measurements of $\psi(\eta)$ with four different values of $\eta=\eta_{1},\eta_{2},\eta_{3},\eta_{4}$, we compute the following quantities,
$\Delta\psi_{1}$ $\equiv\psi(\eta_{1})-\psi(\eta_{2})=\alpha_{-}\left(A\cos 2\theta-B\sin 2\theta\right)$ (14)
$\Delta\psi_{2}$ $\equiv\psi(\eta_{3})-\psi(\eta_{4})=\alpha_{-}\left(C\cos 2\theta-D\sin 2\theta\right),$ (15)
where,
$A$ $=\cos 2\eta_{1}-\cos 2\eta_{2}$ (16)
$B$ $=\sin 2\eta_{1}-\sin 2\eta_{2}$ (17)
$C$ $=\cos 2\eta_{3}-\cos 2\eta_{4}$ (18)
$D$ $=\sin 2\eta_{3}-\sin 2\eta_{4}.$ (19)
We then take the ratio of $\Delta\psi_{1}$ and $\Delta\psi_{2}$,
$r$ $\equiv\frac{\Delta\psi_{1}}{\Delta\psi_{2}}=\frac{A\cos 2\theta-B\sin 2\theta}{C\cos 2\theta-D\sin 2\theta}=\frac{A-B\tan 2\theta}{C-D\tan 2\theta}.$ (20)
Solving the above for $\theta$,
$\theta=\frac{1}{2}\arctan\left(\frac{A-rC}{B-rD}\right).$ (21)
Once the value of $\theta$ is obtained, we can use Equation 12 to compute the value of $\alpha_{-}$ as,
$\alpha_{-}=\arctan\left[\frac{\tan\left\{\psi(\eta)\right\}}{\cos(2\theta+2\eta)}\right].$ (22)
III Birefringence characterization of KAGRA sapphire ITMs
Sapphire is a uniaxial crystal with optical axis called c-axis. KAGRA’s sapphire test masses are fabricated so that the input surface is perpendicular to the c-axis. Light propagating in the test mass substrate will travel along the c-axis and there should be no birefringence in theory. However, local impurities and inhomogeneities in the crystal, and also mechanical and thermal stress changes (e.g., due to the suspension system supporting the mirror) can cause depolarization. These local defects will distort the polarization distribution of the light field in the transverse plane. From Equation 2 we can see the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00. In this case, there will be no depolarization at your input polarization. However, gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates which is determined by $\alpha_{+}$ in Equation 2. To achieve this requirement and to minimize the influence of crystal inhomogeneity, ion beam figuring technique is used during the polishing process in which the thickness of the substrate is precisely adjusted over the surface to make sure the TWE rms is within a few nanometers [18]. This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimized simultaneously. The wavefront error and depolarization is a trade off in birefringent medium.
III.1 TWE of KAGRA sapphire test mass
Figure 3: TWE maps of KAGRA’s current ITMY within the 140 mm aperture measured with linearly polarized light with different $\theta$.
The sapphire ITMs have been installed in KAGRA and directly birefringence mapping is impossible. Characterization of the current sapphire test masses has been studied in [31]. TWE maps of ITMs were measured several times with linearly polarized light by rotating the sample at different angles, which is equivalent with measurements by rotating the input polarization. TWE maps were supposed to be measured at ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$ and ${}^{\circ}$. Methods described in Section II have a rigid requirement on the angular control. However, the measurements in [31] were not optimized for birefringence characterization. So, the rotation angle were not precisely controlled at that time. By comparing locations of the markers on the edge of the mirror in different maps, we were able to recover the real rotation angle (shown in Figure 3). One should notice that TWEs of KAGRA current ITMs measured with linearly polarized light are almost one-order of magnitude larger than specifications, because the vendor used circularly polarized light for ion beam figuring during polishing process. It can be seen from Figure 3, the ${}^{\circ}$ map and the ${}^{\circ}$ map have almost opposite values (also works for ${}^{\circ}$ and ${}^{\circ}$ maps). If summing two maps together, one can derive the TWE for circularly polarized light, which has very small residuals.
III.2 Constructing birefringence maps from TWE
The biggest challenge for birefringence characterizations of KAGRA’s current ITMs comes from the fact that previously TWE measurement setup was not optimized for birefringence study. According to Equation 20 in Section II.3, in order to extract birefringence information, we need subtractions of two pairs of TWE maps and then take the division. In typical Fizeau interferometer measurements, low order effects like the piston and tilt of the surface are not measured. When doing measurements with different $\theta$, the distance and alignment between the sample and reference mirror have changed after rotating the sample. This means that the subtraction of two maps will contain unknown piston and tilt. The numerator and denominator in Equation 20 containing these unknown terms will give wrong birefringence information.
The center of the map also changes after rotating the sample. But we found the decentering is very small (less than 2 mm) and its effect can be negligible when compared with effects from piston and tilt. In addition, we found astigmatisms in original TWE maps. The left plot in Figure 4 shows a TWE map after removing the curvature. A clear astigmatism term can be seen from the map, which is plotted in the right. We found the rotation of this astigmatism does not change when rotating the sample. So, we think the astigmatism comes from the gravity induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer. The astigmatism is removed for each TWE map individually.
Figure 4: Raw TWE map containing astigmatism which mainly comes from deformation due to gravity. As it is not exactly in the vertical direction, we think the surface of the reference sphere used in the Fizeau interferometer is slightly ellipsoidal. After removing the astigmatism, the TWE is shown in the top left plot in Figure 3.
From Equation 21, $\theta$ is wrapped in the range of ($-45^{\circ}$, $45^{\circ}$) due to $\arctan$ function and we need to unwrap the data. Phase unwrapping techniques are widely used in optical metrology and interferometry [32, 33]. In our case, the problem of unknown piston and tilt is that it will give wrong $\theta$ and cause failure of unwrapping. There are algorithms improving phase unwrapping robustness again noise [34, 35, 36, 37]. Most of these methods require the noise to be small compared with the phase change. But the unmeasured piston and tilt in our TWE map can be large. As we take the division of two subtractions, the error will be further amplified when the denominator in Equation 20 is close to zero. The left plot in Figure 5 shows our initially derived $\theta$. The unwrapping algorithm cannot work properly for such kind of map (shown in the right). This results in that the unwrapped surface is discontinuous. The algorithm used in this paper is according to [38, 39].
Figure 5: The left plot shows initially derived $\theta$ map according to Equation 20. Unknown piston and tilt will give wrong $\theta$ and cause failure of unwrapping shown in the right where the unwrapped surface is discontinuous. When performing parameter search for pistons, the 2D matrix is expended to 1D vector along the arrow direction. Each arrow represents a row in the 477$\times$477 data matrix.
In order to recover the correct $\theta$, we perform parameter search for piston terms in the numerator and denominator in Equation 20 (while tilt terms are removed individually for each TWE map and ignored here). This yields
$r=\frac{\psi(\eta_{1})-\psi(\eta_{2})+p_{1}}{\psi(\eta_{3})-\psi(\eta_{4})+p_{2}},$ (23)
where $p_{1}$ and $p_{2}$ are the missing piston terms. For each pair of $p_{1}$ and $p_{2}$ values, we unwrap the derived raw $\theta$ map. Here, we assume the unmeasured pistons are the main cause for the failure of unwrapping. The two-dimensional unwrapped map is then expended to one-dimensional data along the arrow direction shown in the right in Figure 5. If the unwrapped map is a continuous surface, the one-dimensional data should be a continuous curve. So the overall continuity factor of the map can be simply described by
$\mathcal{C}=\sum_{k=1}^{n}(\epsilon_{k+1}-\epsilon_{k})^{2},$ (24)
where $\epsilon_{k}$ is the TWE value at each point of the one-dimensional data and $n$ is the total number of the data. Lower value of $\mathcal{C}$ means a smoother surface while higher value of $\mathcal{C}$ means more value jumps over the surface (see the right plot in Figure 5). Continuity factors of ITMX and ITMY are shown in the right in Figure 6. For each pair of $p_{1}$ and $p_{2}$, $\theta$ and $\alpha_{-}$ can be derived. Then we calculate the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ in Equation 2. This gives the field of linearly polarized light in perpendicular direction with the input polarization. It’s beam shape is shown in the left in Figure 6 with different piston values.
Figure 6: Parameter scan for piston term for ITMX (top) and ITMY (bottom). Plots in the left are single-bounce reflected (passing through the mirror substrate twice) beam shapes in p-polarization (the input is pure s-polarization) with different pistons applied to subtractions of TWEs. Numbers in white are the ratio percentage of beam power in p-polarization over the total input power. Plots in the right are continuity factors of $\theta$ map after unwrapping. It can be seen that $p_{1}$ mainly influence the continuity factor while $p_{2}$ determines the beam shape and power. We search for the correct pistons by eye where continuity factors are small and beam shapes match with measurements shown in Figure 8. We think the beam shape shown in red square gives the best estimation as a consideration of continuity factor and measured beam shape. But actually beams in the green area all have similar shapes with power varying a lot.
III.3 Comparison with in-situ measurements
We have performed birefringence measurements for the sapphire ITMs in KAGRA. A simplified schematic of KAGRA configuration is shown in Figure 7. The laser after the Faraday Isolator is in pure s-polarization. The Power Recycling Mirror (PRM), Signal Recycling Mirror (SRM) and End Test Masses (ETMX and ETMY) are all misaligned so that the beam is reflected back only by ITMs. The plot shows the birefringence measurement for ITMX where ITMY is misaligned. Coatings of the beam splitter (BS) are optimized for s-polarization with a 50%:50% ratio for reflection and transmission at ${}^{\circ}$ incidence. The ratio for p-polarization is 20%:80% according to the coating design. Since there are no direct measurements for p-polarization, practical reflection and transmission may be different from design values. The reflected beams from ITMs are measured at POP and POS ports. There are photodetectors (PD) and cameras for both s-polarization and p-polarization at POP, while there are only photodetectors at POS.
Figure 7: Simplified schematic of KAGRA configuration.
Table 1: Comparison between the measured beam power in p-polarization at POP and POS ports with calculations from $\theta$ and $\alpha_{-}$ maps. Numbers are percentages of the p-polarization power among total power reflected from ITMX or ITMY.
POP POS Calculation
ITMX $(3.65\pm 0.02)\%$ $(4.98\pm 0.01)\%$ 2.5%
ITMY $(8.07\pm 0.04)\%$ $(11.13\pm 0.02)\%$ 6.7%
The measured beam power in p-polarization (as a percentage of the total reflected power) is shown in Table 1, where the effect of unbalanced reflection and transmission of the BS for p-polarization is calibrated (according to coating design values). In principle, the measured beam power content in p-polarization should be the same for POP and POS ports. But in our measurements they are different from each other. We think one reason could be that the practical reflection and transmission of the BS for p-polarization deviate from design. Another reason could be due to the clipping and ghost beam. When the measurement was done, gate valves were added to vacuum tubes whose diameters are just slightly bigger than the beam size. We also found there were a lot of ghost and scattering beams at POP and POS. Though we dumped most of them, there may be still some left and seen by the photodetectors.
Measured beam shapes are shown in the right in Figure 8. By comparing our calculated beam shapes from the previous section with practical measurements, we can find the correct pair of $p_{1}$ and $p_{2}$ by eye which gives both low continuity factor and good agreement with measured beam shapes (see Figure 6). The beam power in p-polarization are then calculated (values are shown in Table 1) with the most suitable pair of $p_{1}$ and $p_{2}$ (marked with red square in Figure 6). The estimated beam power from our maps is smaller than measurements. We think one of the reason is that the unknown tilt term in the measured TWE maps will influence the beam power. By adding tilt influences in the parameter scan, it is possible to increase the calculated power in p-polarization while keeping the beam shape unchanged (or changed by little). In this paper, we don’t perform the full parameter scan for tilt, since the purpose of this paper is to show the feasibility of the method of extracting birefringence information from TWE measurements. An example of the influence of tilt is presented in the Appendix LABEL:app:_tilt.
Figure 8: Comparison between calculated p-polarization beam shape and KAGRA measurements. Dashed circles in white represent the Gaussian beam diameter.
Another factor that can affect our estimated power is the beam spot position on the mirror. The beam spots on ITMs in KAGRA are not exactly in the center but a couple of centimeters away where the power recycling gain and arm transmission power are higher, which is probably due to the homogeneous birefringence effect. Figure 9 shows our estimation how much the beam power in p-polarization will change when the beam is miscentered.
Figure 9: Power of p-polarization when the beam spot is moved away from the mirror center. The x-axis represents the direction of the decenter and $r$ is the distance between the beam spot center and mirror center. Birefringence effect gets smaller when moving the beam towards southeast on ITMX and towards north on ITMY, which can also be seen from Figure 8.
The derived $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY are shown in Figure 10. From these maps, we can see the birefringence in KAGRA’s current ITMs are quite inhomogeneous. If we compare these maps with the beam shape in p-polarization shown in Figure 8, it can be found the beam shape is mainly determined by the shape of $\theta$ map. This is because $\alpha_{-}$ is small in our case. The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarization from Equation 7.
Figure 10: Calibrated $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY. Dashed circles represent the laser beam diameter on mirrors. The plots shown here are the one-way differential phase $\alpha_{-}/2$. The unit of $\alpha_{-}$ is nanometer and the laser wavelength 1064 nm corresponds to a radian of $2\pi$.
As discussed above, a good knowledge of piston and tilt in TWE measurements, as well as the location of the beam spot are necessary in order to precisely estimate the birefringence effect. For more accurate characterizations, the previous setup requires improvements and optimizations for birefringence study.
IV Conclusion and discussion
We have established a method to reconstruct the 2D birefringence information of current KAGRA’s sapphire ITMs from TWE measurements. By combining several TWE measurements with a Fizeau interferometer using linearly polarization in different rotation angles, we are able to successfully construct the birefringence map over a 200 mm area in diameter. Our estimated beam shape for single bounce reflected field in orthogonal polarization shows a good match with in-situ measurements in KAGRA.
It is worth to point out that these TWE measurements were not intended for birefringence characterizations. More accurate reconstruction requires improvements of the previous setup which should target for birefringence study. One possible improvement could be enabling the control of the polarization rotation angle of the input laser in the Fizeau interferometer rather than rotating the tested mirror, as this will remove the influence of the low order (piston, tilt and astigmatism caused by deformation due to gravity) error term in TWE. For future gravitational wave detectors, since most potential materials feasible for cryogenic operation are birefringent, birefringence has become an important parameter for their core optics. This technique of birefringence mapping described in this paper is fast and simple. It can be crucial for future detectors where larger test masses are used. In addition, specifications of mirror birefringence requires a full understanding of how the birefringence affects the performance of the interferometer. The derived birefringence map in this paper can be used to study the effects of inhomogeneous birefringence in future detectors and help us to gain insights of this knowledge.
Acknowledgements.
We would like to thank Miyoki Shinji, Akutsu Tomotada, Takafumi Ushiba and Masahide Tamaki in KAGRA for their support and help for the measurements. We also thank Masaki Ando, Kentaro Somiya, Kentaro Komori and Satoru Takano for insightful discussions. This work was supported by JSPS KAKENHI Grant No. JP20H05854, No. 23H01205, and by JST PRESTO Grant No. JPMJPR200B. KAGRA is supported by MEXT, JSPS Leading-edge Research Infrastructure Program, JSPS Grant-in-Aid for Specially Promoted Research 26000005, JSPS Grant-inAid for Scientific Research on Innovative Areas 2905: JP17H06358, JP17H06361 and JP17H06364, JSPS Core-to-Core Program A. Advanced Research Networks, JSPS Grantin-Aid for Scientific Research (S) 17H06133 and 20H05639 , JSPS Grant-in-Aid for Transformative Research Areas (A) 20A203: JP20H05854, the joint research program of the Institute for Cosmic Ray Research, University of Tokyo, National Research Foundation (NRF), Computing Infrastructure Project of Global Science experimental Data hub Center (GSDC) at KISTI, Korea Astronomy and Space Science Institute (KASI), and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS), AS Grid Center (ASGC) and the National Science and Technology Council (NSTC) in Taiwan under grants including the Rising Star Program and Science Vanguard Research Program, Advanced Technology Center (ATC) of NAOJ, and Mechanical Engineering Center of KEK.
Appendix A Effective Jones matrix of a non-uniform medium along the optical axis
In Section II, we assumed that the transmission of light through a substrate can be described by a single Jones matrix (Equation 2). At a glance, this seems to imply that we also assume the birefringence properties ($\theta$, $n_{e}$ and $n_{o}$) to be constant along the depth of the substrate. However, this assumption does not hold for most real substrates. For the sapphire substrates of the KAGRA mirrors, for example, the fluctuation of birefringence is supposed to come from local mis-alignment of crystal axes, possibly caused by residual internal stress, lattice defects and so on. Then, the birefringence properties must vary not only laterally but also along the depth of the substrate. Then a question arises: can we use a single Jones matrix to describe a complex birefringent medium? In this section, we will address this question.
Figure 11: Splitting a substrate into thin slices.
We start from splitting a thick substrate into thin slices as shown in Figure 11. Each slice is thin enough that the birefringence properties can be regarded as constant within it. Transmission of light through such a thin slice can be described by a single Jones matrix of the form in Equation 2.
As we make the slice infinitesimally thin, $\alpha_{-}$ becomes also infinitesimally small. From Equation 2, we can expand the Jones matrix with $\alpha_{-}$ and leave only its first order terms (the common phase term $\alpha_{+}$ is omitted here),
$M$ $=\begin{pmatrix}\cos\alpha_{-}+i\cos 2\theta\sin\alpha_{-}&-i\sin 2\theta\sin\alpha_{-}\\ -i\sin 2\theta\sin\alpha_{-}&\cos\alpha_{-}-i\cos 2\theta\sin\alpha_{-}\end{pmatrix}$ (27)
$\sim\begin{pmatrix}1+i\cos 2\theta\cdot\alpha_{-}&-i\sin 2\theta\cdot\alpha_{-}\\ -i\sin 2\theta\cdot\alpha_{-}&1-i\cos 2\theta\cdot\alpha_{-}\end{pmatrix}.$ (30)
The next step is to consider the transmission of light through two consecutive slices having different Jones matrices:
$M_{1}$ $=\begin{pmatrix}1+i\cos 2\theta_{1}\cdot\alpha_{1}&-i\sin 2\theta_{1}\cdot\alpha_{1}\\ -i\sin 2\theta_{1}\cdot\alpha_{1}&1-i\cos 2\theta_{1}\cdot\alpha_{1}\end{pmatrix},$ (33)
$M_{2}$ $=\begin{pmatrix}1+i\cos 2\theta_{2}\cdot\alpha_{2}&-i\sin 2\theta_{2}\cdot\alpha_{2}\\ -i\sin 2\theta_{2}\cdot\alpha_{2}&1-i\cos 2\theta_{2}\cdot\alpha_{2}\end{pmatrix}.$ (36)
Such a process can be described by the multiplication of the two Jones matrices,
$M_{12}=M_{2}\cdot M_{1}=\begin{pmatrix}1+i\left(\cos 2\theta_{1}\cdot\alpha_{1}+\cos 2\theta_{2}\cdot\alpha_{2}\right)&-i\left(\sin 2\theta_{1}\cdot\alpha_{1}+\sin 2\theta_{2}\cdot\alpha_{2}\right)\\ -i\left(\sin 2\theta_{1}\cdot\alpha_{1}+\sin 2\theta_{2}\cdot\alpha_{2}\right)&1-i\left(\cos 2\theta_{1}\cdot\alpha_{1}+\cos 2\theta_{2}\cdot\alpha_{2}\right)\end{pmatrix}.$ (39)
By comparing Equation 39 and Equation 30, we need to find $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ which satisfy the following conditions,
$\alpha_{\mathrm{e}}\cos 2\theta_{\mathrm{e}}$ $=\cos 2\theta_{1}\cdot\alpha_{1}+\cos 2\theta_{2}\cdot\alpha_{2}$ (40)
$\alpha_{\mathrm{e}}\sin 2\theta_{\mathrm{e}}$ $=\sin 2\theta_{1}\cdot\alpha_{1}+\sin 2\theta_{2}\cdot\alpha_{2}.$ (41)
Finding such a combination of $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ is always possible. Then we can express $M_{12}$ using these parameters as,
$M_{12}=\begin{pmatrix}1+i\cos 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}&-i\sin 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}\\ -i\sin 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}&1-i\cos 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}\end{pmatrix}.$ (42)
This is an effective Jones matrix for the transmission of two consecutive thin slices. By repeating this process, we can always find a single Jones matrix of the form shown in Equation 2 representing the transmission of a thicker substrate. Therefore, the use of Equation 2 for the analysis of an inhomogeneous thick substrate is justified.
Appendix B Influence of tilt of TWE maps on the estimation of beam shape and beam power
In Section III.2, we show that the piston, which is usually ignored in an interferometric measurement with Fizeau interferometers, however, is an important factor in birefringence measurements with the method described in this paper. The tilt, also ignored in TWE measurements, has a similar effect. In this section, we give an example of its influence to our birefringence characterizations.
The tilt of a surface can be described by a combination of the first-order Zernike polynomials $Z_{1}^{-1}$ and $Z_{1}^{1}$ (shwon in Figure 12). Thus we need to add two degrees of freedom of tilt for nominator and denominator respectively in Equation 23 and this ends up with four more degrees of freedom ($a$, $b$, $c$ and $d$) for parameter scan, which yields
$r=\frac{\psi(\eta_{1})-\psi(\eta_{2})+p_{1}+aZ_{1}^{-1}+bZ_{1}^{1}}{\psi(\eta_{3})-\psi(\eta_{4})+p_{2}+cZ_{1}^{-1}+dZ_{1}^{1}}.$ (43)
Figure 12: The first order coefficients of Zernike polynomials. Any tilt of a surface can be described with a combination of them with different amplitudes.
According to Figure 6, the change in the nominator in Equation 43 mainly influences the continuity factor, while the change in the denominator has a larger effect on the beam shape and power. Here, we set $a=b=0$ to see the influence of tilt on beam shape and power. Pistons $p_{1}$ and $p_{2}$ are fixed according to values giving results in red squares in Figure 6. Figure 13 shows the change of beam shape and power when we apply different amount of tilt to the denominator. The results labeled in red squares show a good match with birefringence measurements at POP in terms of both beam shape and beam power in p-polarization.
Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarization with different magnitude of tilt.

### v2

@@
-Characterization of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements
+Characterisation of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements
 Abstract
-Future gravitational wave detectors are going to cool down their test masses to cryogenic temperatures to reduce thermal noise.
-Crystalline materials are considered most promising for making test masses and their coatings due to their excellent thermal and optical properties at low temperatures.
-However, birefringence due to local impurities and inhomogeneities in the crystal can degrade the performance of the detector.
-Birefringence measurement, or birefringence mapping over a two-dimensional area has become important.
-This paper describes a method of fast birefringence measurement for a large sample by simply combining a series of transmission wavefront error measurements using linearly polarized light with Fizeau interferometers.
-With this method, the birefringence inhomogeneity of two KAGRA’s input test masses with diameter of 22 cm is fully reconstructed.
-The birefringence information is then used to calculate the transverse beam shape of the light fields in orthogonal polarization directions when passing through the substrate.
-The calculated beam shapes show a good match with in-situ measurements using KAGRA interferometer.
-This technique is crucial for birefringence characterization of test masses in future detectors where even larger size are used.
+Cooling down test masses to cryogenic temperatures is a way to reduce the thermal noise of gravitational wave detectors.
+Crystalline materials are considered the most promising materials for fabricating cryogenic test masses and their coatings because of their excellent thermal and optical properties at low temperatures.
+However, birefringence owing to local impurities and inhomogeneities in the crystal can degrade the performance of the detector.
+The birefringence measurement or mapping over a two-dimensional area is thus important.
+This study describes a method for fast birefringence measurements of a large sample by simply combining a series of transmission wavefront error measurements using linearly polarised light with Fizeau interferometers.
+Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses with a diameter of 22 cm was fully reconstructed.
+The birefringence information was then used to calculate the transverse beam shapes of the light fields in orthogonal polarisation directions when passing through the substrate.
+The calculated beam shapes showed a good agreement with the in-situ measurements using the KAGRA interferometer.
+This technique is crucial for birefringence characterisation of test masses in future detectors, where even larger sizes are used.
 I Introduction
-Thermal noise is one of the main noise sources for future gravitational wave detectors [1, 2].
-To reduce thermal noise, future detectors like LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures.
-Cosmic Explorer [6] also considers cryogenic operation for an upgrade.
-KAGRA [7, 8, 9, 10] is the only one operating at cryogenic temperatures and pioneering this technique among current large scale ground-based detectors.
-All these detectors apply large size high-quality mirrors as their test masses forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves.
-Fused silica is commonly used to make test masses at room temperature, but it is not compatible with low temperatures.
-Crystalline materials are considered most promising to make cryogenic test masses.
-Sapphire was selected as the material for KAGRA’s test masses [11] due to its excellent thermal and optical properties at low temperature [12, 13].
-However, recent studies reveal that birefringence of crystalline substrates and coatings can degrade the performance of the detector [14, 15, 16, 17, 18, 19].
+Thermal noise is one of the main noise sources for gravitational wave detectors [1, 2].
+To reduce thermal noise, future detectors such as the LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures.
+Cosmic Explorer [6] considers cryogenic operations for upgrades.
+KAGRA [7, 8, 9, 10] is the only detector operating at cryogenic temperatures and pioneers this technique among the current large-scale ground-based detectors.
+All these detectors use large, high-quality mirrors as their test masses, forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves.
+Fused silica is commonly used to make test masses at room temperature; however, it is not compatible with low temperatures.
+Crystalline materials are considered to be the most promising materials for fabricating cryogenic test masses.
+Sapphire was selected as the material for KAGRA’s test masses [11] because of its excellent thermal and optical properties at low temperatures [12, 13].
+However, recent studies have revealed that the birefringence of crystalline substrates and coatings can degrade detector performance [14, 15, 16, 17, 18, 19].
 In this case, the birefringence fluctuates in magnitude and direction throughout the substrate of the mirror.
-Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on the optical performance.
-There are many studies on two-dimensional birefringence measurements for transparent materials [14, 15, 20, 21, 22, 23, 24].
-These measurements are either using a single-beam scanning technique, or expanding the beam to a large aperture and imaging with a sensor array.
-The scanning method is rather time-consuming for large size samples, e.g., mirrors with diameters usually more than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [25, 26].
-The phase-shifting method in digital interferometry has been previously proposed for fast 2D birefringence measurements with a very high accuracy [27, 28, 29].
-Fizeau interferometers are one of the most accurate methods to extract the phase information of an optical component (reflection phase or transmission phase) over a large area.
-Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test.
-In the presence of birefringence, however, a simple TWE measurement does not reveal the entire information about the substrate.
-In this paper we describe a method of combining a series of TWE map measurements using linearly polarized light in an appropriate way to fully reconstruct the refractive index information including birefringence.
-With this method, the birefringence inhomogeneity of two KAGRA’s input test masses (ITM) is extracted.
-This birefringence information is then used to calculate the transverse beam shape of the light fields in orthogonal polarization directions when passing through the substrate.
-The calculated beam shapes show a good match with in-situ beam shape measurements.
+Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on optical performance.
+There are several studies on two-dimensional birefringence measurements of transparent materials [14, 15, 20, 21, 22, 23, 24].
+These measurements use either a single-beam scanning technique or expand the beam to a large aperture for imaging using a sensor array.
+The scanning method is rather time-consuming for large sample sizes, for example, mirrors with diameters usually greater than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [25, 26].
+A phase-shifting method in digital interferometry has been proposed for fast 2D birefringence measurements with high accuracy [27, 28, 29].
+Fizeau interferometry is among the most accurate methods for extracting the phase information of an optical component (reflection or transmission phase) over a large area.
+The Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test.
+However, in the presence of birefringence a simple TWE measurement does not reveal all the information about the substrate.
+In this study, we describe a method for combining a series of TWE map measurements using linearly polarised light in an appropriate way to fully reconstruct refractive index information, including birefringence.
+Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses (ITM) is extracted.
+This birefringence information was then used to calculate the transverse beam shape of the light fields in orthogonal polarisation directions when passing through the substrate.
+The calculated beam shapes were in good agreement with the in-situ beam shape measurements.
 The method of extracting birefringence information from TWE measurements with Fizeau interferometers is introduced in Section II.
-Birefringence characterization of KAGRA’s current ITMs is described in Section III.
-Our conclusions are summarized in Section IV.
+The birefringence characterisation of KAGRA’s current ITMs is described in Section III, and the conclusions are summarised in Section IV.
 II Generating a birefringence map with a Fizeau interferometer
-Figure 1 shows a schematic of a Fizeau interferometer.
-We analyze the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD.
-By comparing the results with and without the sample inserted into the cavity formed by the reference planes 1 and 2, we can obtain the information about the optical thickness of the sample.
-Since the optical thickness is the physical thickness times the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflecting coating) and S2 (surface with highly-reflective coating) precisely and subtracting out the variation of physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
 Figure 1: Schematic of TWE measurement with a Fizeau interferometer.
-S1 is the surface with anti-reflecting coating and S2 is the surface with highly-reflective coating.
-The interference between light fields $E_{\mathrm{1}}$ and $E_{\mathrm{2}}$ reflected by reference planes 1 and 2 respectively gives the TWE information of the sample.
-In order to extract the birefringence information, measurements have to be performed using a linearly polarized light.
-To understand the behavior of such measurements, let us first consider what happens to a linearly polarized light going through a birefringent substrate.
-II.1 Transmission of a linearly polarized light through a birefringent medium
-Let us concentrate on an optical axis going through a single column of the substrate as shown in Figure 2.
-We define our preferred polarization axes called s and p in the vertical and horizontal directions with respect to the laboratory frame.
-The extra-ordinary and ordinary axes of the crystal in that particular column is assumed to be rotated by an angle $\theta$ with respect to our polarization axes.
-The input beam to the Fizeau interferometer is purely polarized in s ($E_{\mathrm{s}}$).
-Figure 2: Transmission of a linearly polarized optical field through a column of a sample substrate.
-We will use the Jones matrix formalism to analyze the conversion of the polarization state [30]. Here, the polarization state of light is represented by a vector,
+S1 is the surface with anti-reflection coating and S2 is the surface with highly reflective coating.
+The interference between light fields $E_{\mathrm{1}}$ and $E_{\mathrm{2}}$ reflected by reference planes 1 and 2, respectively, obtains the TWE information of the sample.
+Figure 1 shows a schematic diagram of the Fizeau interferometer.
+We analyse the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD.
+By comparing the results obtained with and without the sample inserted into the cavity formed by reference planes 1 and 2, we obtained information on the optical thickness of the sample.
+Because the optical thickness is the physical thickness multiplied by the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflection coating) and S2 (surface with highly reflective coating) precisely and subtracting the variation in the physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
+To extract birefringence information, measurements must be performed using linearly polarised light.
+To understand the behaviour of such measurements, we first considered the behaviour of linearly polarised light passing through a birefringent substrate.
+II.1 Transmission of a linearly polarised light through a birefringent medium
+Consider optical axis passing through a single substrate column, as shown in Figure 2.
+We defined our preferred polarisation axes, s and p, in the vertical and horizontal directions with respect to the laboratory frame.
+The extraordinary and ordinary axes of the crystal in that particular column are assumed to be rotated by an angle $\theta$ with respect to our polarisation axes.
+The input beam to the Fizeau interferometer was purely polarised in s ($E_{\mathrm{s}}$).
+Figure 2: Transmission of a linearly polarised optical field through a column of a sample substrate.
+The Jones matrix formalism was used to analyse the conversion of the polarisation state [30]. The polarisation state of light is represented by the following vector:
 $\vec{V}=\begin{pmatrix}E_{\mathrm{s}}\\ E_{\mathrm{p}}\end{pmatrix},$ (1)
-where $E_{\mathrm{s}}$ and $E_{\mathrm{p}}$ are the complex amplitudes of s- and p-polarized electric fields.
-After passing through the substrate twice (a round trip through the substrate), the polarization state is converted to
+where $E_{\mathrm{s}}$ and $E_{\mathrm{p}}$ are the complex amplitudes of the s- and p-polarised electric fields, respectively.
+After passing through the substrate twice (a round trip through the substrate), the polarisation state was converted to:
 $\vec{V}^{\prime}=M\cdot\vec{V},\quad\mathrm{with}\quad M=e^{i\alpha_{+}}\begin{pmatrix}e^{i\alpha_{-}}\cos^{2}\theta+e^{-i\alpha_{-}}\sin^{2}\theta&-i\sin\alpha_{-}\sin 2\theta\\ -i\sin\alpha_{-}\sin 2\theta&e^{-i\alpha_{-}}\cos^{2}\theta+e^{i\alpha_{-}}\sin^{2}\theta\end{pmatrix},$ (2)
-where $M$ is the round-trip Jones matrix of the substrate.
-We denote the indices of refraction for the extra-ordinary and ordinary axes as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively.
-Then we can compute the round trip phase changes of optical fields polarized in those axes:
+where $M$ denotes the round-trip Jones matrix of the substrate.
+The indices of refraction for the extraordinary and ordinary axes are denoted as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively.
+Then, we can compute the round- trip phase changes of the optical fields polarised along the axes:
 $\alpha_{\mathrm{e}}\equiv 2\pi\frac{2dn_{\mathrm{e}}}{\lambda},$ (3)
 $\alpha_{\mathrm{o}}\equiv 2\pi\frac{2dn_{\mathrm{o}}}{\lambda},$ (4)
-where $d$ is the thickness of the substrate, $\lambda$ is the wavelength of the light.
-Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes,
+where $d$ is the thickness of the substrate, and $\lambda$ is the wavelength of the light.
+Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes as follows:
 $\alpha_{+}\equiv\frac{\alpha_{\mathrm{e}}+\alpha_{\mathrm{o}}}{2},$ (5)
@@
 In most cases, the common phase is not important and we will ignore $\alpha_{+}$ in the following discussion. Note that the Jones matrix also works for a non-uniform medium (discussed in Appendix A).
-For a purely s-polarized input field, $\vec{V}=(1,0)^{\mathrm{T}}$, the field returning from the substrate becomes,
+For a purely s-polarised input field, $\vec{V}=(1,0)^{\mathrm{T}}$, the field returning from the substrate becomes,
 $\vec{V}^{\prime}=\begin{pmatrix}E_{\mathrm{s}}\\ E_{\mathrm{p}}\end{pmatrix}=\begin{pmatrix}e^{i\alpha_{-}}\cos^{2}\theta+e^{-i\alpha_{-}}\sin^{2}\theta\\ -i\sin\alpha_{-}\sin 2\theta\end{pmatrix}.$ (7)
-This $\vec{V}^{\prime}$ is the same as the field $E_{2}$ in Figure 1 at the CCD, except for the common phase term, which is ignored.
+This $\vec{V}^{\prime}$ is the same as the field $E_{2}$, as shown in Figure 1, at the CCD, except for the common phase term, which is ignored.
 II.2 Interference with the reference field
 The field $\vec{V}^{\prime}$ computed above is superposed with the reference field ($E_{1}$ in Figure 1).
-The reference field $\vec{V}_{\mathrm{r}}$ is purely polarized in s and has a certain phase offset $\phi$ from $\vec{V}^{\prime}$,
+The reference field $\vec{V}_{\mathrm{r}}$ is purely polarised in s and has a certain phase offset $\phi$ from $\vec{V}^{\prime}$,
 $\vec{V}_{\mathrm{r}}=\begin{pmatrix}E_{\mathrm{r}}\\ 0\end{pmatrix}=\begin{pmatrix}e^{i\phi}\\ 0\end{pmatrix}.$ (8)
-The optical power detected by the CCD after the interference of $\vec{V}^{\prime}$ and $\vec{V}_{\mathrm{r}}$ is,
+The optical power detected by the CCD after the interference of $\vec{V}^{\prime}$ and $\vec{V}_{\mathrm{r}}$ is as follows:
 $P=(\vec{V}_{\mathrm{r}}+\vec{V}^{\prime})\cdot(\vec{V}_{\mathrm{r}}+\vec{V}^{\prime})^{*}.$ (9)
-By plugging Equations 7 and 8 into Equation 9, we get,
+By substituting Equations (7) and (8) into Equation (9), we obtain:
 $P$ $=|E_{\mathrm{s}}+E_{\mathrm{r}}|^{2}+|E_{\mathrm{p}}|^{2}$
 $=|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}+|E_{\mathrm{r}}|^{2}+E_{\mathrm{s}}E_{\mathrm{r}}^{*}+E_{\mathrm{s}}^{*}E_{\mathrm{r}}.$ (10)
-From the conservation of energy, $|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}=1$.
-From Equation 8, $|E_{\mathrm{r}}|^{2}=1$.
-We will ignore those constant terms in the following discussion.
-Then what is interesting to us is,
+Based on energy conservation, $|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}=1$. From Equation (8), $|E_{\mathrm{r}}|^{2}=1$. We ignore these constant terms in the following discussion. However, note that
 $\Delta P$ $=E_{\mathrm{s}}E_{\mathrm{r}}^{*}+E_{\mathrm{s}}^{*}E_{\mathrm{r}}$
 $=2\cos^{2}\theta\cos(\phi-\alpha_{-})+2\sin^{2}\theta\cos(\phi+\alpha_{-}).$ (11)
-If the s-polarization axis is aligned with the extra-ordinary axis ($\theta=0$), $\Delta P=2\cos(\phi-\alpha_{-})$.
-Usually, a Fizeau interferometer is equipped with a mechanism to change the position of the reference plane 1, for example, by mounting the reference flat on a piezo stage.
-Therefore, we can scan the relative phase $\phi$ by known amounts and fit the measured variation of $\Delta P$ to know the value of $\alpha_{-}$.
-In the general case of $\theta\neq 0$, the Fizeau measurement obtains the phase of the field $E_{\mathrm{s}}$, i.e.
-$\psi=\arg(E_{\mathrm{s}})$.
+If the s-polarisation axis is aligned with the extra-ordinary axis ($\theta=0$), $\Delta P=2\cos(\phi-\alpha_{-})$.
+A Fizeau interferometer is equipped with a mechanism to change the position of the reference plane 1, for example, by mounting a reference flat on a piezo stage.
+Therefore, we can scan the relative phase $\phi$ by known amounts and fit the measured variation of $\Delta P$ to obtain the value of $\alpha_{-}$.
+In the general case of $\theta\neq 0$, the Fizeau measurement yields the phase of the field $E_{\mathrm{s}}$, that is, $\psi=\arg(E_{\mathrm{s}})$.
 II.3 Extraction of $\alpha_{-}$ and $\theta$ from several TWE maps
-When the orientation of the extra-ordinary axis is not known, which is almost always the case, we need to take several TWE maps with different orientation of the sample or the measurement beam polarization.
+When the orientation of the extraordinary axis is unknown, which is almost always the case, we must obtain several TWE maps with different orientations of the sample or measurement beam polarisation.
 Then combine them to extract $\theta$ and $\alpha_{-}$.
 As explained in the previous section, a TWE measurement by a Fizeau interferometer yields a transmission phase $\psi$ at each point on the sample surface.
-Let us assume that we rotate the sample by an angle $\eta$, or equivalently the input polarization orientation by $-\eta$.
+Let us assume that we rotate the sample by an angle $\eta$, or equivalently the input polarisation orientation by $-\eta$.
 This is equivalent to changing the orientation of the extra-ordinary axis from $\theta$ to $\theta+\eta$.
-Let us denote the measured phase with an $\eta$ rotation of the sample as $\psi(\eta)$. From Equation 7,
+Let us denote the measured phase with an $\eta$ rotation of the sample as $\psi(\eta)$. From Equation (7),
 $\psi(\eta)$ $=\arg\left[E_{\mathrm{s}}(\theta+\eta)\right]$
@@
 We can see $\psi(\eta)$ as a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape.
-Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation 12 to obtain the best fit values of $\theta$ and $\alpha_{-}$.
-If $\alpha_{-}$ can be assumed to be very small, which is the case for KAGRA mirrors, we can expand Equation 12 with $\alpha_{-}$, leaving only the $\mathcal{O}(\alpha_{-})$ terms, as
+Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation (12) to obtain the best fit values of $\theta$ and $\alpha_{-}$.
+If $\alpha_{-}$ can be assumed to be very small, which is the case for KAGRA mirrors, we can expand Equation (12) with $\alpha_{-}$, leaving only the $\mathcal{O}(\alpha_{-})$ terms, as follows:
 $\psi(\eta)$ $\sim\cos(2\theta+2\eta)\alpha_{-}$
@@
 $D$ $=\sin 2\eta_{3}-\sin 2\eta_{4}.$ (19)
-We then take the ratio of $\Delta\psi_{1}$ and $\Delta\psi_{2}$,
+We then use the ratio of $\Delta\psi_{1}$ and $\Delta\psi_{2}$,
 $r$ $\equiv\frac{\Delta\psi_{1}}{\Delta\psi_{2}}=\frac{A\cos 2\theta-B\sin 2\theta}{C\cos 2\theta-D\sin 2\theta}=\frac{A-B\tan 2\theta}{C-D\tan 2\theta}.$ (20)
@@
 $\theta=\frac{1}{2}\arctan\left(\frac{A-rC}{B-rD}\right).$ (21)
-Once the value of $\theta$ is obtained, we can use Equation 12 to compute the value of $\alpha_{-}$ as,
+Once the value of $\theta$ is obtained, we can use Equation (12) to compute the value of $\alpha_{-}$ as follows:
 $\alpha_{-}=\arctan\left[\frac{\tan\left\{\psi(\eta)\right\}}{\cos(2\theta+2\eta)}\right].$ (22)
-III Birefringence characterization of KAGRA sapphire ITMs
-Sapphire is a uniaxial crystal with optical axis called c-axis.
-KAGRA’s sapphire test masses are fabricated so that the input surface is perpendicular to the c-axis.
-Light propagating in the test mass substrate will travel along the c-axis and there should be no birefringence in theory.
-However, local impurities and inhomogeneities in the crystal, and also mechanical and thermal stress changes (e.g., due to the suspension system supporting the mirror) can cause depolarization.
-These local defects will distort the polarization distribution of the light field in the transverse plane.
-From Equation 2 we can see the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00.
-In this case, there will be no depolarization at your input polarization.
-However, gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates which is determined by $\alpha_{+}$ in Equation 2.
-To achieve this requirement and to minimize the influence of crystal inhomogeneity, ion beam figuring technique is used during the polishing process in which the thickness of the substrate is precisely adjusted over the surface to make sure the TWE rms is within a few nanometers [18].
-This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimized simultaneously.
-The wavefront error and depolarization is a trade off in birefringent medium.
+III Birefringence characterisation of KAGRA sapphire ITMs
+Table 1: Rms values of TWE specifications for the KAGRA test masses and measurements with a circular/linear polarisation beam within an aperture of 140 mm [31].
+Circular Linear
+Specification polarisation polarisation
+ITMX $<$5 nm 4.09 nm 25.90 nm
+ITMY $<$5 nm 4.61 nm 32.72 nm
+Sapphire is a uniaxial crystal with an optical axis called the c-axis.
+KAGRA’s sapphire test masses are fabricated such that the input surface is perpendicular to the c-axis.
+Light propagating in the test mass substrate travels along the c-axis, and theoretically, should experience no birefringence.
+However, local impurities and inhomogeneities in the crystal, as well as mechanical and thermal stress (e.g.
+due to the suspension system supporting the mirror), can cause depolarisation.
+These local defects distorted the polarisation distribution of the light field in the transverse plane.
+From Equation (2), we can see that the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00.
+In this case, there was no depolarisation for the input polarisation.
+Gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates.
+To achieve this requirement and minimise the influence of crystal inhomogeneity, when the birefringence was ignored, an ion beam figuring technique was used during the polishing process, in which the thickness of the substrate was precisely adjusted over the surface to ensure that the TWE rms was within a few nanometres [18].
+However, in a birefringent medium, $\alpha_{+}$ and $\alpha_{-}$ in Equation (2) are both functions of thickness.
+This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimised simultaneously by ajusting the thickness of the substrate.
+Wavefront errors and depolarisation are trade-offs in a birefringent medium.
 III.1 TWE of KAGRA sapphire test mass
-Figure 3: TWE maps of KAGRA’s current ITMY within the 140 mm aperture measured with linearly polarized light with different $\theta$.
-The sapphire ITMs have been installed in KAGRA and directly birefringence mapping is impossible.
-Characterization of the current sapphire test masses has been studied in [31].
-TWE maps of ITMs were measured several times with linearly polarized light by rotating the sample at different angles, which is equivalent with measurements by rotating the input polarization.
-TWE maps were supposed to be measured at ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$ and ${}^{\circ}$.
-Methods described in Section II have a rigid requirement on the angular control.
-However, the measurements in [31] were not optimized for birefringence characterization.
-So, the rotation angle were not precisely controlled at that time.
-By comparing locations of the markers on the edge of the mirror in different maps, we were able to recover the real rotation angle (shown in Figure 3).
-One should notice that TWEs of KAGRA current ITMs measured with linearly polarized light are almost one-order of magnitude larger than specifications, because the vendor used circularly polarized light for ion beam figuring during polishing process.
-It can be seen from Figure 3, the ${}^{\circ}$ map and the ${}^{\circ}$ map have almost opposite values (also works for ${}^{\circ}$ and ${}^{\circ}$ maps).
-If summing two maps together, one can derive the TWE for circularly polarized light, which has very small residuals.
+The sapphire ITMs have already been installed in KAGRA, and direct birefringence mapping is impossible now.
+However, characterisation of the sapphire test masses was studied in [31].
+The TWE maps of the ITMs were measured several times with linearly polarised light by rotating the sample at different angles, which was equivalent to the measurements obtained by rotating the input polarisation.
+The TWE maps were measured at the angles of 0, 45, 90 and 135.
+The methods described in Section II require a rigid angular control.
+However, the main motivation of the measurements in [31] was to see the stress induced change in $\alpha_{+}$, not $\alpha_{-}$, by different mounting orientations.
+Therefore, the rotation angle was not precisely controlled.
+To mitigate the problem, we compared the locations of the markers on the edge of the mirrors in different maps.
+In this way, we were able to recover the real rotation angle (Figure 3).
+The TWEs of the KAGRA current ITMs measured with linearly polarised light are around 5 to 6 times larger than the specifications because the vendor used circularly polarised light to measure the path length difference and calculate the removal from surface S2 during the ion beam figuring process (rms values are listed in Table 1).
+As shown in Figure 3, the 0 map and the 88 map have almost opposite values (the same is true for 49 and 130 maps).
+By summing the two maps, we can derive the TWE for circularly polarised light, which has small residuals.
+Figure 3: TWE maps of KAGRA’s current ITMY within the 140 mm aperture measured with linearly polarised light with different $\theta$.
 III.2 Constructing birefringence maps from TWE
-The biggest challenge for birefringence characterizations of KAGRA’s current ITMs comes from the fact that previously TWE measurement setup was not optimized for birefringence study.
-According to Equation 20 in Section II.3, in order to extract birefringence information, we need subtractions of two pairs of TWE maps and then take the division.
-In typical Fizeau interferometer measurements, low order effects like the piston and tilt of the surface are not measured.
+The biggest challenge for the birefringence characterisation of KAGRA’s current ITMs is that the previous TWE measurement setup was not optimised for birefringence studies.
+According to Equation (20) in Section II.3, to extract the birefringence information, we need to subtract two pairs of TWE maps and then divide.
+In typical Fizeau interferometer measurements, low-order effects such as the piston and surface tilt are removed from the results.
 When doing measurements with different $\theta$, the distance and alignment between the sample and reference mirror have changed after rotating the sample.
-This means that the subtraction of two maps will contain unknown piston and tilt.
-The numerator and denominator in Equation 20 containing these unknown terms will give wrong birefringence information.
-The center of the map also changes after rotating the sample.
-But we found the decentering is very small (less than 2 mm) and its effect can be negligible when compared with effects from piston and tilt.
-In addition, we found astigmatisms in original TWE maps.
-The left plot in Figure 4 shows a TWE map after removing the curvature.
-A clear astigmatism term can be seen from the map, which is plotted in the right.
-We found the rotation of this astigmatism does not change when rotating the sample.
-So, we think the astigmatism comes from the gravity induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer.
-The astigmatism is removed for each TWE map individually.
-Figure 4: Raw TWE map containing astigmatism which mainly comes from deformation due to gravity.
-As it is not exactly in the vertical direction, we think the surface of the reference sphere used in the Fizeau interferometer is slightly ellipsoidal.
+This implies that the difference between the two maps contains an unknown piston and tilt.
+The numerator and denominator in Equation (20), which contain these unknown terms, provide incorrect birefringence information.
+The centre of the map also changed after the sample was rotated.
+However, we observed that the decentring was very small (less than 2 mm), and its effect was negligible compared with the effects of the piston and tilt.
+In addition, astigmatism was observed on the original TWE maps.
+The left plot in Figure 4 shows the TWE map after removing the curvature.
+A clear astigmatism term can be observed on the map plotted on the right.
+We observed that the rotation of this astigmatism did not change when the samples were rotated.
+Therefore, we believe that the astigmatism originates from the gravity-induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer.
+Astigmatism was individually removed from each TWE map.
+Figure 4: Raw TWE map containing astigmatism which mainly occurs due to deformation due to gravity.
+As it is not in the vertical direction, we assume that the surface of the reference sphere used in the Fizeau interferometer is slightly ellipsoidal.
 After removing the astigmatism, the TWE is shown in the top left plot in Figure 3.
-From Equation 21, $\theta$ is wrapped in the range of ($-45^{\circ}$, $45^{\circ}$) due to $\arctan$ function and we need to unwrap the data.
+The algorithm used in processing the maps was based on Simtools [32].
+From Equation (21), $\theta$ is wrapped in the range of ($-45^{\circ}$, $45^{\circ}$) due to the $\arctan$ function and we need to unwrap the data.
 Phase unwrapping techniques are widely used in optical metrology and interferometry [33, 34].
-In our case, the problem of unknown piston and tilt is that it will give wrong $\theta$ and cause failure of unwrapping.
-There are algorithms improving phase unwrapping robustness again noise [34, 35, 36, 37].
+In our case, the problem of the unknown piston and tilt is that it will give the wrong $\theta$ and cause failure of unwrapping.
+Some algorithms improve phase unwrapping robustness against noise [35, 36, 37, 38].
 Most of these methods require the noise to be small compared with the phase change.
-But the unmeasured piston and tilt in our TWE map can be large.
-As we take the division of two subtractions, the error will be further amplified when the denominator in Equation 20 is close to zero.
+However, the unmeasured piston and tilt on the TWE map can be large.
+As we take the division of two subtractions, the error will be further amplified when the denominator in Equation (20) is close to zero.
 The left plot in Figure 5 shows our initially derived $\theta$.
-The unwrapping algorithm cannot work properly for such kind of map (shown in the right).
-This results in that the unwrapped surface is discontinuous.
-The algorithm used in this paper is according to [38, 39].
-Figure 5: The left plot shows initially derived $\theta$ map according to Equation 20.
-Unknown piston and tilt will give wrong $\theta$ and cause failure of unwrapping shown in the right where the unwrapped surface is discontinuous.
-When performing parameter search for pistons, the 2D matrix is expended to 1D vector along the arrow direction.
+The unwrapping algorithm does not function properly for such maps (shown on the right).
+This resulted in a discontinuous, unwrapped surface.
+The algorithm used in this study was based on [39, 40].
+Figure 5: The left plot shows initially derived $\theta$ map according to Equation (20).
+The unknown piston and tilt will give wrong $\theta$ and cause failure of unwrapping shown on the right where the unwrapped surface is discontinuous.
+When performing parameter search for pistons, the 2D matrix is expanded to 1D vector along the arrow direction.
 Each arrow represents a row in the 477$\times$477 data matrix.
-In order to recover the correct $\theta$, we perform parameter search for piston terms in the numerator and denominator in Equation 20 (while tilt terms are removed individually for each TWE map and ignored here).
+To recover the correct $\theta$, we perform parameter search for piston terms in the numerator and denominator shown in Equation (20) Tilt terms are removed individually for each TWE map (by removing first-order Zernike polynomials) and ignored here.
 This yields
 $r=\frac{\psi(\eta_{1})-\psi(\eta_{2})+p_{1}}{\psi(\eta_{3})-\psi(\eta_{4})+p_{2}},$ (23)
-where $p_{1}$ and $p_{2}$ are the missing piston terms.
-For each pair of $p_{1}$ and $p_{2}$ values, we unwrap the derived raw $\theta$ map.
-Here, we assume the unmeasured pistons are the main cause for the failure of unwrapping.
-The two-dimensional unwrapped map is then expended to one-dimensional data along the arrow direction shown in the right in Figure 5.
+where $p_{1}$ and $p_{2}$ are missing piston terms.
+For each pair of $p_{1}$ and $p_{2}$ values, the derived raw $\theta$ map was unwrapped.
+Here, we assume that unmeasured pistons are the main cause of unwrapping failure.
+The two-dimensional unwrapped map is then expanded to one-dimensional data along the arrow direction, as shown on the right in Figure 5.
 If the unwrapped map is a continuous surface, the one-dimensional data should be a continuous curve.
-So the overall continuity factor of the map can be simply described by
+Therefore, the overall continuity factor of the map can be simply described by
 $\mathcal{C}=\sum_{k=1}^{n}(\epsilon_{k+1}-\epsilon_{k})^{2},$ (24)
 where $\epsilon_{k}$ is the TWE value at each point of the one-dimensional data and $n$ is the total number of the data.
-Lower value of $\mathcal{C}$ means a smoother surface while higher value of $\mathcal{C}$ means more value jumps over the surface (see the right plot in Figure 5).
-Continuity factors of ITMX and ITMY are shown in the right in Figure 6.
+A lower value of $\mathcal{C}$ indicates a smoother surface, whereas a higher value of $\mathcal{C}$ means that more values jump over the surface (refer the right plot shown in Figure 5).
+The continuity factors of ITMX and ITMY are shown on the right side of Figure 6.
 For each pair of $p_{1}$ and $p_{2}$, $\theta$ and $\alpha_{-}$ can be derived.
-Then we calculate the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ in Equation 2.
-This gives the field of linearly polarized light in perpendicular direction with the input polarization.
-It’s beam shape is shown in the left in Figure 6 with different piston values.
+Subsequently, we calculate the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ in Equation (2).
+This provided a field of linearly polarised light in the direction perpendicular to the input polarisation.
+The beam shape is shown on the left in Figure 6 with different piston values.
 Figure 6: Parameter scan for piston term for ITMX (top) and ITMY (bottom).
-Plots in the left are single-bounce reflected (passing through the mirror substrate twice) beam shapes in p-polarization (the input is pure s-polarization) with different pistons applied to subtractions of TWEs.
-Numbers in white are the ratio percentage of beam power in p-polarization over the total input power.
-Plots in the right are continuity factors of $\theta$ map after unwrapping.
-It can be seen that $p_{1}$ mainly influence the continuity factor while $p_{2}$ determines the beam shape and power.
+Plots on the left are single-bounce reflected (passing through the mirror substrate twice) beam shapes in p-polarisation (the input is pure s-polarisation) with different pistons applied to subtractions of TWEs.
+Numbers in white are the ratio percentage of beam power in p-polarisation over the total input power.
+Plots on the right are continuity factors of $\theta$ map after unwrapping.
+It can be observed that $p_{1}$ mainly influences the continuity factor, whereas $p_{2}$ determines the beam shape and power.
 We search for the correct pistons by eye where continuity factors are small and beam shapes match with measurements shown in Figure 8.
-We think the beam shape shown in red square gives the best estimation as a consideration of continuity factor and measured beam shape.
-But actually beams in the green area all have similar shapes with power varying a lot.
+We assume that the beam shape shown in the red square gives the best estimation as a consideration of the continuity factor and measured beam shape.
+However, beams in the green area all have similar shapes with significantly varying power.
 III.3 Comparison with in-situ measurements
-We have performed birefringence measurements for the sapphire ITMs in KAGRA.
-A simplified schematic of KAGRA configuration is shown in Figure 7.
-The laser after the Faraday Isolator is in pure s-polarization.
-The Power Recycling Mirror (PRM), Signal Recycling Mirror (SRM) and End Test Masses (ETMX and ETMY) are all misaligned so that the beam is reflected back only by ITMs.
-The plot shows the birefringence measurement for ITMX where ITMY is misaligned.
-Coatings of the beam splitter (BS) are optimized for s-polarization with a 50%:50% ratio for reflection and transmission at ${}^{\circ}$ incidence.
-The ratio for p-polarization is 20%:80% according to the coating design.
-Since there are no direct measurements for p-polarization, practical reflection and transmission may be different from design values.
-The reflected beams from ITMs are measured at POP and POS ports.
-There are photodetectors (PD) and cameras for both s-polarization and p-polarization at POP, while there are only photodetectors at POS.
+Birefringence measurements were performed for the sapphire ITMs at KAGRA.
+A simplified schematic of the KAGRA configuration used during this measurement is shown in Figure 7.
+The laser after the Faraday Isolator was in pure s-polarisation.
+The power recycling mirror (PRM), signal recycling mirror (SRM), and end test masses (ETMX and ETMY) are all misaligned such that the beam is reflected only by the ITMs.
+The figure shows the particular setup for measuring the birefringence of ITMX, where ITMY is misaligned.
+The coatings of the beam splitter (BS) were optimised for s-polarisation with a 50%:50% ratio for reflection and transmission at 45 incidence.
+The p-polarisation ratio was 20%:80% according to the coating design.
+Because there is no direct measurement of coating reflectivity for p-polarisation, the practical reflection and transmission may be different from the design values.
+The beams reflected from the ITM were measured at the POP and POS ports.
+Photodetectors (PD) and cameras are used for both s-polarisation and p-polarisation at the POP, whereas only photodetectors are used at the POS.
 Figure 7: Simplified schematic of KAGRA configuration.
-Table 1: Comparison between the measured beam power in p-polarization at POP and POS ports with calculations from $\theta$ and $\alpha_{-}$ maps.
-Numbers are percentages of the p-polarization power among total power reflected from ITMX or ITMY.
+Table 2: Comparison between the measured beam power in p-polarisation at POP and POS ports with calculations from $\theta$ and $\alpha_{-}$ maps.
+Numbers are percentages of the p-polarisation power with respect to the total power reflected from the ITMX or ITMY.
 POP POS Calculation
@@
 ITMY $(8.07\pm 0.04)\%$ $(11.13\pm 0.02)\%$ 6.7%
-The measured beam power in p-polarization (as a percentage of the total reflected power) is shown in Table 1, where the effect of unbalanced reflection and transmission of the BS for p-polarization is calibrated (according to coating design values).
-In principle, the measured beam power content in p-polarization should be the same for POP and POS ports.
-But in our measurements they are different from each other.
-We think one reason could be that the practical reflection and transmission of the BS for p-polarization deviate from design.
-Another reason could be due to the clipping and ghost beam.
-When the measurement was done, gate valves were added to vacuum tubes whose diameters are just slightly bigger than the beam size.
-We also found there were a lot of ghost and scattering beams at POP and POS.
-Though we dumped most of them, there may be still some left and seen by the photodetectors.
-Measured beam shapes are shown in the right in Figure 8.
-By comparing our calculated beam shapes from the previous section with practical measurements, we can find the correct pair of $p_{1}$ and $p_{2}$ by eye which gives both low continuity factor and good agreement with measured beam shapes (see Figure 6).
-The beam power in p-polarization are then calculated (values are shown in Table 1) with the most suitable pair of $p_{1}$ and $p_{2}$ (marked with red square in Figure 6).
-The estimated beam power from our maps is smaller than measurements.
-We think one of the reason is that the unknown tilt term in the measured TWE maps will influence the beam power.
-By adding tilt influences in the parameter scan, it is possible to increase the calculated power in p-polarization while keeping the beam shape unchanged (or changed by little).
-In this paper, we don’t perform the full parameter scan for tilt, since the purpose of this paper is to show the feasibility of the method of extracting birefringence information from TWE measurements.
-An example of the influence of tilt is presented in the Appendix LABEL:app:_tilt.
-Figure 8: Comparison between calculated p-polarization beam shape and KAGRA measurements. Dashed circles in white represent the Gaussian beam diameter.
-Another factor that can affect our estimated power is the beam spot position on the mirror.
-The beam spots on ITMs in KAGRA are not exactly in the center but a couple of centimeters away where the power recycling gain and arm transmission power are higher, which is probably due to the homogeneous birefringence effect.
-Figure 9 shows our estimation how much the beam power in p-polarization will change when the beam is miscentered.
-Figure 9: Power of p-polarization when the beam spot is moved away from the mirror center.
-The x-axis represents the direction of the decenter and $r$ is the distance between the beam spot center and mirror center.
-Birefringence effect gets smaller when moving the beam towards southeast on ITMX and towards north on ITMY, which can also be seen from Figure 8.
+The measured beam power in p-polarisation (as a percentage of the total reflected power) is listed in Table 2, where the effect of the unbalanced reflection and transmission of the BS on the p-polarisation is calibrated (according to the coating design values).
+In principle, the measured beam power content in p-polarisation should be the same for the POP and POS ports.
+However, our measurements are different from each other.
+One reason for this could be that the practical reflection and transmission of the BS for p-polarisation deviate from the design.
+Another reason could be clipping and ghost beams.
+When the measurement was performed, gate valves with diameters slightly larger than the beam diameter were added to the vacuum tubes.
+We also found many ghost and scattering beams at the POP and POS.
+Although we dumped most of them, there may be still some remaining and are seen by the photodetectors.
+The measured beam shapes are shown on the right side in Figure 8.
+By comparing our calculated beam shapes from the previous section with practical measurements, we can find the correct pair of $p_{1}$ and $p_{2}$ visually, which gives both a low continuity factor and good agreement with the measured beam shapes (refer Figure 6).
+The beam power in p-polarisation was then calculated (the values are listed in Table 2) using the most suitable pair of $p_{1}$ and $p_{2}$ (marked with the red square shown in Figure 6).
+The beam power estimated from our maps was lower than that from the measurements.
+One reason is that the unknown tilt term in the measured TWE maps influences the beam power.
+By adding tilt influences in the parameter scan, it is possible to increase the calculated power in p-polarisation while maintaining the beam shape unchanged (or slightly changed).
+In this study, we did not perform a full parameter scan for tilt because the purpose was to demonstrate the feasibility of the method of extracting birefringence information from TWE measurements.
+An example of the influence of the tilt is presented in Appendix B.
+Figure 8: Comparison between calculated p-polarisation beam shape and KAGRA measurements.
+Dashed circles in white represent the Gaussian beam diameter, the width at which the amplitude is $1/e$ of its value in the centre.
+Another factor that can affect the estimated power is the position of the beam spot on the mirror.
+The beam spots on the ITMs in KAGRA are not exactly in the centre but are a couple of centimetres away where the power recycling gain and arm transmission power are higher, which is probably due to the inhomogeneous birefringence effect.
+Figure 9 shows our estimate of the extent to which the p-polarised beam power changes when the beam is miscentered.
+Figure 9: Power of p-polarisation when the beam spot is moved away from the mirror centre.
+The x-axis represents the direction of the decentre and r is the distance between the beam spot centre and mirror centre.
+Birefringence effect gets smaller when moving the beam towards southeast on ITMX and towards north on ITMY, as shown in Figure 8.
 The derived $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY are shown in Figure 10.
-From these maps, we can see the birefringence in KAGRA’s current ITMs are quite inhomogeneous.
-If we compare these maps with the beam shape in p-polarization shown in Figure 8, it can be found the beam shape is mainly determined by the shape of $\theta$ map.
+From these maps, we can see that the birefringence in KAGRA’s current ITMs is inhomogeneous.
+If we compare these maps with the beam shape in p-polarisation shown in Figure 8, it can be found that the beam shape is mainly determined by the shape of $\theta$ map.
 This is because $\alpha_{-}$ is small in our case.
-The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarization from Equation 7.
+The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarisation from Equation (7).
 Figure 10: Calibrated $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY.
@@
 The plots shown here are the one-way differential phase $\alpha_{-}/2$.
-The unit of $\alpha_{-}$ is nanometer and the laser wavelength 1064 nm corresponds to a radian of $2\pi$.
-As discussed above, a good knowledge of piston and tilt in TWE measurements, as well as the location of the beam spot are necessary in order to precisely estimate the birefringence effect.
-For more accurate characterizations, the previous setup requires improvements and optimizations for birefringence study.
+The unit of $\alpha_{-}$ is nanometre and the laser wavelength 1064 nm corresponds to a radian of $2\pi$.
+As discussed above, a good knowledge of the piston and tilt in TWE measurements and the location of the beam spot is necessary to precisely estimate the birefringence effect.
+For more accurate characterisation, the Fizeau interferometer requires improvements and optimisations for the birefringence study.
 IV Conclusion and discussion
-We have established a method to reconstruct the 2D birefringence information of current KAGRA’s sapphire ITMs from TWE measurements.
-By combining several TWE measurements with a Fizeau interferometer using linearly polarization in different rotation angles, we are able to successfully construct the birefringence map over a 200 mm area in diameter.
-Our estimated beam shape for single bounce reflected field in orthogonal polarization shows a good match with in-situ measurements in KAGRA.
-It is worth to point out that these TWE measurements were not intended for birefringence characterizations.
-More accurate reconstruction requires improvements of the previous setup which should target for birefringence study.
-One possible improvement could be enabling the control of the polarization rotation angle of the input laser in the Fizeau interferometer rather than rotating the tested mirror, as this will remove the influence of the low order (piston, tilt and astigmatism caused by deformation due to gravity) error term in TWE.
-For future gravitational wave detectors, since most potential materials feasible for cryogenic operation are birefringent, birefringence has become an important parameter for their core optics.
-This technique of birefringence mapping described in this paper is fast and simple.
-It can be crucial for future detectors where larger test masses are used.
-In addition, specifications of mirror birefringence requires a full understanding of how the birefringence affects the performance of the interferometer.
-The derived birefringence map in this paper can be used to study the effects of inhomogeneous birefringence in future detectors and help us to gain insights of this knowledge.
+We established a method for reconstructing the 2D birefringence information of current KAGRA sapphire ITMs from TWE measurements.
+By combining several TWE measurements with a Fizeau interferometer using linear polarisation at different rotation angles, we successfully constructed a birefringence map over a 200 mm area in diameter.
+Our estimated beam shape for a single-bounce reflected field in orthogonal polarisation shows a good match with the in situ measurements in KAGRA.
+Note that these TWE measurements were not intended for birefringence characterisations.
+More accurate reconstruction requires improvements to the previous setup, which should target birefringence studies.
+One possible improvement could be enabling control of the polarisation rotation angle of the input laser in the Fizeau interferometer rather than rotating the tested mirror, as this will remove the influence of the low-order (piston, tilt, and astigmatism caused by deformation due to gravity) error term in the TWE.
+For future gravitational wave detectors, birefringence has become an important parameter for their core optics because most potential materials that are suitable for cryogenic operation are birefringent.
+The birefringence-mapping technique described in this study is fast and simple.
+This is crucial for future detectors with larger test masses.
+In addition, the specifications of mirror birefringence require a complete understanding of how birefringence affects the interferometer performance.
+The derived birefringence map in this study can be used to study the effects of inhomogeneous birefringence in future detectors and help us gain insights into this knowledge.
 Acknowledgements.
-We would like to thank Miyoki Shinji, Akutsu Tomotada, Takafumi Ushiba and Masahide Tamaki in KAGRA for their support and help for the measurements.
+We would like to thank Shinji Miyoki and Tomotada Akutsu at KAGRA for their support for the measurements.
 We also thank Masaki Ando, Kentaro Somiya, Kentaro Komori and Satoru Takano for insightful discussions.
@@
 JP20H05854, No.
+22H01228, No.
 23H01205, and by JST PRESTO Grant No.
@@
 Appendix A Effective Jones matrix of a non-uniform medium along the optical axis
-In Section II, we assumed that the transmission of light through a substrate can be described by a single Jones matrix (Equation 2).
-At a glance, this seems to imply that we also assume the birefringence properties ($\theta$, $n_{e}$ and $n_{o}$) to be constant along the depth of the substrate.
+In Section II, we assumed that the transmission of light through a substrate can be described by a single Jones matrix (Equation (2)).
+This seems to imply that we also assume the birefringence properties ($\theta$, $n_{e}$ and $n_{o}$) to be constant along the depth of the substrate.
 However, this assumption does not hold for most real substrates.
-For the sapphire substrates of the KAGRA mirrors, for example, the fluctuation of birefringence is supposed to come from local mis-alignment of crystal axes, possibly caused by residual internal stress, lattice defects and so on.
-Then, the birefringence properties must vary not only laterally but also along the depth of the substrate.
-Then a question arises: can we use a single Jones matrix to describe a complex birefringent medium?
-In this section, we will address this question.
+For the sapphire substrates of KAGRA mirrors, for example, the fluctuation of birefringence is supposed to originate from the local misalignment of crystal axes, possibly caused by residual internal stress and lattice defects.
+Therefore, the birefringence properties must vary not only laterally but also along the depth of the substrate.
+Thus, the question arises: can we use a single Jones matrix to describe a complex birefringent medium?
+This section addresses this question.
 Figure 11: Splitting a substrate into thin slices.
-We start from splitting a thick substrate into thin slices as shown in Figure 11.
-Each slice is thin enough that the birefringence properties can be regarded as constant within it.
-Transmission of light through such a thin slice can be described by a single Jones matrix of the form in Equation 2.
+We began by splitting a thick substrate into thin slices, as shown in Figure 11.
+Each slice is sufficiently thin such that the birefringence properties can be regarded as constant.
+The transmission of light through such a thin slice can be described by a single Jones matrix, as shown in Equation (2).
 As we make the slice infinitesimally thin, $\alpha_{-}$ becomes also infinitesimally small.
-From Equation 2, we can expand the Jones matrix with $\alpha_{-}$ and leave only its first order terms (the common phase term $\alpha_{+}$ is omitted here),
+From Equation (2), we can expand the Jones matrix with $\alpha_{-}$ and leave only its first order terms (the common phase term $\alpha_{+}$ is omitted here),
 $M$ $=\begin{pmatrix}\cos\alpha_{-}+i\cos 2\theta\sin\alpha_{-}&-i\sin 2\theta\sin\alpha_{-}\\ -i\sin 2\theta\sin\alpha_{-}&\cos\alpha_{-}-i\cos 2\theta\sin\alpha_{-}\end{pmatrix}$ (25)
 $\sim\begin{pmatrix}1+i\cos 2\theta\cdot\alpha_{-}&-i\sin 2\theta\cdot\alpha_{-}\\ -i\sin 2\theta\cdot\alpha_{-}&1-i\cos 2\theta\cdot\alpha_{-}\end{pmatrix}.$ (26)
-The next step is to consider the transmission of light through two consecutive slices having different Jones matrices:
+Next, we considered the transmission of light through two consecutive slices with different Jones matrices:
 $M_{1}$ $=\begin{pmatrix}1+i\cos 2\theta_{1}\cdot\alpha_{1}&-i\sin 2\theta_{1}\cdot\alpha_{1}\\ -i\sin 2\theta_{1}\cdot\alpha_{1}&1-i\cos 2\theta_{1}\cdot\alpha_{1}\end{pmatrix},$ (27)
@@
 $M_{12}=M_{2}\cdot M_{1}=\begin{pmatrix}1+i\left(\cos 2\theta_{1}\cdot\alpha_{1}+\cos 2\theta_{2}\cdot\alpha_{2}\right)&-i\left(\sin 2\theta_{1}\cdot\alpha_{1}+\sin 2\theta_{2}\cdot\alpha_{2}\right)\\ -i\left(\sin 2\theta_{1}\cdot\alpha_{1}+\sin 2\theta_{2}\cdot\alpha_{2}\right)&1-i\left(\cos 2\theta_{1}\cdot\alpha_{1}+\cos 2\theta_{2}\cdot\alpha_{2}\right)\end{pmatrix}.$ (29)
-By comparing Equation 39 and Equation 30, we need to find $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ which satisfy the following conditions,
+By comparing Equation (29) and Equation (26), we find $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$, which satisfy the following conditions:
 $\alpha_{\mathrm{e}}\cos 2\theta_{\mathrm{e}}$ $=\cos 2\theta_{1}\cdot\alpha_{1}+\cos 2\theta_{2}\cdot\alpha_{2}$ (30)
 $\alpha_{\mathrm{e}}\sin 2\theta_{\mathrm{e}}$ $=\sin 2\theta_{1}\cdot\alpha_{1}+\sin 2\theta_{2}\cdot\alpha_{2}.$ (31)
-Finding such a combination of $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ is always possible. Then we can express $M_{12}$ using these parameters as,
+Finding such a combination of $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ is always possible. Then we express $M_{12}$ using these parameters as follows:
 $M_{12}=\begin{pmatrix}1+i\cos 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}&-i\sin 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}\\ -i\sin 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}&1-i\cos 2\theta_{\mathrm{e}}\cdot\alpha_{\mathrm{e}}\end{pmatrix}.$ (32)
-This is an effective Jones matrix for the transmission of two consecutive thin slices.
-By repeating this process, we can always find a single Jones matrix of the form shown in Equation 2 representing the transmission of a thicker substrate.
-Therefore, the use of Equation 2 for the analysis of an inhomogeneous thick substrate is justified.
+This is an effective Jones matrix for transmitting two consecutive thin slices.
+By repeating this process, we can obtain a single Jones matrix of the form shown in Equation (2), which represents the transmission of a thick substrate.
+Therefore, the use of Equation (2) for the analysis of thick, inhomogeneous substrates is justified.
 Appendix B Influence of tilt of TWE maps on the estimation of beam shape and beam power
-In Section III.2, we show that the piston, which is usually ignored in an interferometric measurement with Fizeau interferometers, however, is an important factor in birefringence measurements with the method described in this paper.
-The tilt, also ignored in TWE measurements, has a similar effect.
-In this section, we give an example of its influence to our birefringence characterizations.
-The tilt of a surface can be described by a combination of the first-order Zernike polynomials $Z_{1}^{-1}$ and $Z_{1}^{1}$ (shwon in Figure 12).
-Thus we need to add two degrees of freedom of tilt for nominator and denominator respectively in Equation 23 and this ends up with four more degrees of freedom ($a$, $b$, $c$ and $d$) for parameter scan, which yields
+Figure 12: The first order coefficients of Zernike polynomials. Tilts of a surface can be described using a combination of them with different amplitudes.
+Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarisation with different magnitude of tilt.
+In Section III.2, we show that the piston, which is typically ignored in interferometric measurements using Fizeau interferometers, is an important factor in birefringence measurements using the method described in this paper.
+Tilt, which is also ignored in TWE measurements, has a similar effect.
+In this section, we provide an example of its influence on birefringence characterisation.
+The tilt of the surface can be described by a combination of the first-order Zernike polynomials $Z_{1}^{-1}$ and $Z_{1}^{1}$ (shwon in Figure 12).
+Therefore, we need to add two degrees of freedom of tilt for the numerator and denominator in Equation (23), which results in four more degrees of freedom ($a$, $b$, $c$ and $d$) for the parameter scan, yielding
 $r=\frac{\psi(\eta_{1})-\psi(\eta_{2})+p_{1}+aZ_{1}^{-1}+bZ_{1}^{1}}{\psi(\eta_{3})-\psi(\eta_{4})+p_{2}+cZ_{1}^{-1}+dZ_{1}^{1}}.$ (33)
-Figure 12: The first order coefficients of Zernike polynomials. Any tilt of a surface can be described with a combination of them with different amplitudes.
-According to Figure 6, the change in the nominator in Equation 43 mainly influences the continuity factor, while the change in the denominator has a larger effect on the beam shape and power.
-Here, we set $a=b=0$ to see the influence of tilt on beam shape and power.
-Pistons $p_{1}$ and $p_{2}$ are fixed according to values giving results in red squares in Figure 6.
-Figure 13 shows the change of beam shape and power when we apply different amount of tilt to the denominator.
-The results labeled in red squares show a good match with birefringence measurements at POP in terms of both beam shape and beam power in p-polarization.
-Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarization with different magnitude of tilt.
+According to Figure 6, the change in the numerator of Equation (33) mainly influences the continuity factor, whereas the change in the denominator has a larger effect on the beam shape and power.
+Here, we set $a=b=0$ to determine the influence of the tilt on the beam shape and power.
+Pistons $p_{1}$ and $p_{2}$ are fixed according to their values, resulting in the red squares shown in Figure 6.
+Figure 13 shows the change in the beam shape and power when different amounts of tilt are applied to the denominator.
+The results labelled with red squares show a good match with the birefringence measurements at the POP in terms of both the beam shape and beam power in p-polarisation.

### v3

@@
 The birefringence information was then used to calculate the transverse beam shapes of the light fields in orthogonal polarisation directions when passing through the substrate.
-The calculated beam shapes showed a good agreement with the in-situ measurements using the KAGRA interferometer.
+It was possible to find a calculated beam shape consistent with in-situ measurements using the KAGRA interferometer.
 This technique is crucial for birefringence characterisation of test masses in future detectors, where even larger sizes are used.
@@
 Thermal noise is one of the main noise sources for gravitational wave detectors [1, 2].
-To reduce thermal noise, future detectors such as the LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures.
-Cosmic Explorer [6] considers cryogenic operations for upgrades.
+To reduce thermal noise, future detectors such as the Einstein Telescope [3, 4] and LIGO Voyager [5] concept are designed to operate at cryogenic temperatures.
+Cosmic Explorer [6, 7] considers cryogenic operations for upgrades.
 KAGRA [8, 9, 10, 11] is the only detector operating at cryogenic temperatures and pioneers this technique among the current large-scale ground-based detectors.
 All these detectors use large, high-quality mirrors as their test masses, forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves.
-Fused silica is commonly used to make test masses at room temperature; however, it is not compatible with low temperatures.
+Fused silica is commonly used to make test masses at room temperature; however, it is not compatible with cryogenic detectors due to its large mechanical loss at low temperatures.
 Crystalline materials are considered to be the most promising materials for fabricating cryogenic test masses.
 Sapphire was selected as the material for KAGRA’s test masses [12] because of its excellent thermal and optical properties at low temperatures [13, 14].
-However, recent studies have revealed that the birefringence of crystalline substrates and coatings can degrade detector performance [14, 15, 16, 17, 18, 19].
-In this case, the birefringence fluctuates in magnitude and direction throughout the substrate of the mirror.
+However, recent studies have shown that the birefringence of crystalline substrates and coatings can degrade detector performance [15, 16, 17, 18, 19, 20].
+The birefringence fluctuates spatially in magnitude and direction throughout the mirror substrate.
 Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on optical performance.
-There are several studies on two-dimensional birefringence measurements of transparent materials [14, 15, 20, 21, 22, 23, 24].
+There are several studies on two-dimensional birefringence measurements of transparent materials [15, 16, 21, 22, 23, 24, 25, 26].
 These measurements use either a single-beam scanning technique or expand the beam to a large aperture for imaging using a sensor array.
@@
 However, in the presence of birefringence a simple TWE measurement does not reveal all the information about the substrate.
-In this study, we describe a method for combining a series of TWE map measurements using linearly polarised light in an appropriate way to fully reconstruct refractive index information, including birefringence.
-Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses (ITM) is extracted.
-This birefringence information was then used to calculate the transverse beam shape of the light fields in orthogonal polarisation directions when passing through the substrate.
-The calculated beam shapes were in good agreement with the in-situ beam shape measurements.
-The method of extracting birefringence information from TWE measurements with Fizeau interferometers is introduced in Section II.
-The birefringence characterisation of KAGRA’s current ITMs is described in Section III, and the conclusions are summarised in Section IV.
+In this study, we describe a method for combining a series of TWE map measurements using linearly polarised light in an appropriate way to fully reconstruct refractive index information including birefringence, which is introduced in Section II.
+Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses (ITM) is extracted, described in Section III.
+This birefringence information was then used to calculate the transverse beam shape of the light fields in orthogonal polarisation directions when passing through the substrate, which is compared with the in-situ beam shape measurements.
+Conclusions and discussions are summarised in Section IV.
 II Generating a birefringence map with a Fizeau interferometer
@@
 We analyse the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD.
-By comparing the results obtained with and without the sample inserted into the cavity formed by reference planes 1 and 2, we obtained information on the optical thickness of the sample.
+By comparing the interference between reference planes 1 and 2, we obtained information on the optical thickness of the sample.
 Because the optical thickness is the physical thickness multiplied by the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflection coating) and S2 (surface with highly reflective coating) precisely and subtracting the variation in the physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
@@
 where $M$ denotes the round-trip Jones matrix of the substrate.
-The indices of refraction for the extraordinary and ordinary axes are denoted as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively.
+The indices of refraction for the ordinary and extraordinary axes are denoted as $n_{\mathrm{o}}$ and $n^{\prime}_{\mathrm{e}}$ [27], respectively.
 Then, we can compute the round- trip phase changes of the optical fields polarised along the axes:
-$\alpha_{\mathrm{e}}\equiv 2\pi\frac{2dn_{\mathrm{e}}}{\lambda},$ (3)
 $\alpha_{\mathrm{o}}\equiv 2\pi\frac{2dn_{\mathrm{o}}}{\lambda},$ (3)
+$\alpha_{\mathrm{e}}\equiv 2\pi\frac{2dn^{\prime}_{\mathrm{e}}}{\lambda},$ (4)
 where $d$ is the thickness of the substrate, and $\lambda$ is the wavelength of the light.
-Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes as follows:
-$\alpha_{+}\equiv\frac{\alpha_{\mathrm{e}}+\alpha_{\mathrm{o}}}{2},$ (5)
-$\alpha_{-}\equiv\frac{\alpha_{\mathrm{e}}-\alpha_{\mathrm{o}}}{2}.$ (6)
+Using $\alpha_{\mathrm{o}}$ and $\alpha_{\mathrm{e}}$, we define the common and differential phase changes as follows:
+$\alpha_{+}\equiv\frac{\alpha_{\mathrm{o}}+\alpha_{\mathrm{e}}}{2}=\frac{4\pi d}{\lambda}(n_{\mathrm{o}}+n^{\prime}_{\mathrm{e}}),$ (5)
+$\alpha_{-}\equiv\frac{\alpha_{\mathrm{o}}-\alpha_{\mathrm{e}}}{2}=\frac{4\pi d}{\lambda}(n_{\mathrm{o}}-n^{\prime}_{\mathrm{e}}).$ (6)
 In most cases, the common phase is not important and we will ignore $\alpha_{+}$ in the following discussion. Note that the Jones matrix also works for a non-uniform medium (discussed in Appendix A).
@@
 $=\arctan\left[\cos(2\theta+2\eta)\tan\alpha_{-}\right].$ (12)
-We can see $\psi(\eta)$ as a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape.
+We can see $\psi(\eta)$ is a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape.
 Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation (12) to obtain the best fit values of $\theta$ and $\alpha_{-}$.
@@
 $\alpha_{-}=\arctan\left[\frac{\tan\left\{\psi(\eta)\right\}}{\cos(2\theta+2\eta)}\right].$ (22)
-III Birefringence characterisation of KAGRA sapphire ITMs
+III Birefringence characterisation of KAGRA sapphire Input test masses
 Table 1: Rms values of TWE specifications for the KAGRA test masses and measurements with a circular/linear polarisation beam within an aperture of 140 mm [33].
@@
 ITMY $<$5 nm 4.61 nm 32.72 nm
-Sapphire is a uniaxial crystal with an optical axis called the c-axis.
+Sapphire is a uniaxial crystal with an optical axis called the c-axis that has no intrinsic birefringence.
 KAGRA’s sapphire test masses are fabricated such that the input surface is perpendicular to the c-axis.
@@
 However, local impurities and inhomogeneities in the crystal, as well as mechanical and thermal stress (e.g.
-due to the suspension system supporting the mirror), can cause depolarisation.
+due to the suspension system supporting the mirror), can cause birefringence.
 These local defects distorted the polarisation distribution of the light field in the transverse plane.
-From Equation (2), we can see that the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00.
-In this case, there was no depolarisation for the input polarisation.
 Gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates.
-To achieve this requirement and minimise the influence of crystal inhomogeneity, when the birefringence was ignored, an ion beam figuring technique was used during the polishing process, in which the thickness of the substrate was precisely adjusted over the surface to ensure that the TWE rms was within a few nanometres [18].
-However, in a birefringent medium, $\alpha_{+}$ and $\alpha_{-}$ in Equation (2) are both functions of thickness.
-This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimised simultaneously by ajusting the thickness of the substrate.
-Wavefront errors and depolarisation are trade-offs in a birefringent medium.
+To achieve this requirement and minimise the influence of crystal inhomogeneity, an ion beam figuring technique was used during the polishing process, in which the thickness of the substrate was precisely adjusted over the surface to ensure that the TWE rms was within a few nanometres [19].
+However, this technique optimizes the TWE for the beam with a certain polarisation state, and it cannot reduce birefringence.
 III.1 TWE of KAGRA sapphire test mass
@@
 As shown in Figure 3, the 0 map and the 88 map have almost opposite values (the same is true for 49 and 130 maps).
-By summing the two maps, we can derive the TWE for circularly polarised light, which has small residuals.
+By summing the two maps, we can derive the TWE for circularly polarised light, whose rms magnitude is below 5 nm and matches with the specification.
 Figure 3: TWE maps of KAGRA’s current ITMY within the 140 mm aperture measured with linearly polarised light with different $\theta$.
@@
 The biggest challenge for the birefringence characterisation of KAGRA’s current ITMs is that the previous TWE measurement setup was not optimised for birefringence studies.
-According to Equation (20) in Section II.3, to extract the birefringence information, we need to subtract two pairs of TWE maps and then divide.
+According to Equation (20) in Section II.3, to extract the birefringence information, we need to subtract two pairs of TWE maps and then divide one by the other.
 In typical Fizeau interferometer measurements, low-order effects such as the piston and surface tilt are removed from the results.
@@
 The left plot in Figure 4 shows the TWE map after removing the curvature.
-A clear astigmatism term can be observed on the map plotted on the right.
+An astigmatism term can be observed on the map plotted on the right.
 We observed that the rotation of this astigmatism did not change when the samples were rotated.
@@
 From these maps, we can see that the birefringence in KAGRA’s current ITMs is inhomogeneous.
+$\theta$ fluctuates a lot in the range of ±$\pm 50^{\circ}$ due to the fluctuation of the c-axis.
+The one-way differential phase (retardation) $\alpha_{-}/2$ is in the order of 100 nm, which corresponds to $\Delta n=n_{\mathrm{o}}-n^{\prime}_{\mathrm{e}}=6.7\times 10^{-7}$ (with the thickness of the mirror $d=0.15$ m).
 If we compare these maps with the beam shape in p-polarisation shown in Figure 8, it can be found that the beam shape is mainly determined by the shape of $\theta$ map.
@@
 22H01228, No.
-23H01205, and by JST PRESTO Grant No.
+23H01205, No.
+23K19053, No.
+24K00649, and by JST PRESTO Grant No.
 JPMJPR200B.
@@
 This section addresses this question.
-Figure 11: Splitting a substrate into thin slices.
 We began by splitting a thick substrate into thin slices, as shown in Figure 11.
@@
 The transmission of light through such a thin slice can be described by a single Jones matrix, as shown in Equation (2).
+Figure 11: Splitting a substrate into thin slices.
 As we make the slice infinitesimally thin, $\alpha_{-}$ becomes also infinitesimally small.
@@
 Appendix B Influence of tilt of TWE maps on the estimation of beam shape and beam power
-Figure 12: The first order coefficients of Zernike polynomials. Tilts of a surface can be described using a combination of them with different amplitudes.
-Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarisation with different magnitude of tilt.
 In Section III.2, we show that the piston, which is typically ignored in interferometric measurements using Fizeau interferometers, is an important factor in birefringence measurements using the method described in this paper.
@@
 $r=\frac{\psi(\eta_{1})-\psi(\eta_{2})+p_{1}+aZ_{1}^{-1}+bZ_{1}^{1}}{\psi(\eta_{3})-\psi(\eta_{4})+p_{2}+cZ_{1}^{-1}+dZ_{1}^{1}}.$ (33)
+Figure 12: The first order coefficients of Zernike polynomials. Tilts of a surface can be described using a combination of them with different amplitudes.
+Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarisation with different magnitude of tilt.
 According to Figure 6, the change in the numerator of Equation (33) mainly influences the continuity factor, whereas the change in the denominator has a larger effect on the beam shape and power.